Consider a consumer who consumes only two goods, and y. His utility over these two goods is given by U(x,y) = xy. The budget constraint of the consumer is given by 3x + 9y = 216, where 3 is the price of good x, 9 is the price of good y and 216 is the total income of the consumer.

(a) Find the optimal quantities of good and that the consumer is going to consume. Show the solution in a graph. What level of utility is the consumer going to achieve with this bundle?


(b) Now assume that the price of good increases to 6. Find the new optimal consumption bundle and show it in a graph


1
Expert's answer
2022-01-10T09:53:49-0500

Solution:

a.). Optimal quantities of good x and y:

"\frac{MUx}{MUy} = \frac{Px}{Py}"


MUx = "\frac{\partial U} {\partial x} = y"


MUy = "\frac{\partial U} {\partial y} = x"

 

"\frac{y}{x} = \frac{3}{9}"

x = 3y

Substitute in the budget constraint:

216 = 3x + 9y

216 = 3(3y) + 9y = 9y + 9y = 18y

216 = 18y

Y = 12

X = 3y = 3 "\times" 12 = 36

Optimal quantities of x and y = 36, 12

Level of utility = xy = 36 x 12 = 432

The graph is as below:



 

b.). "\frac{y}{x} = \frac{6}{9}"

x = 1.5y

Substitute in the budget constraint:

216 = 6x + 9y

216 = 6(1.5y) + 9y = 9y + 9y = 18y

216 = 18y

Y = 12

X = 1.5y = 1.5 "\times" 12 = 18

Optimal quantities of x and y = 18, 12

Level of utility = xy = 18 "\times" 12 = 216


The graph is as below: